# Generative Context Distillation

### v1

@@ Abstract
Prompts used in recent large language model based applications are often fixed and lengthy, leading to significant computational overhead. To address this challenge, we propose Generative Context Distillation (GCD), a lightweight prompt internalization method that employs a joint training approach. This method not only replicates the behavior of models with prompt inputs but also generates the content of the prompt along with reasons for why the model’s behavior should change accordingly. We demonstrate that our approach effectively internalizes complex prompts across various agent-based application scenarios. For effective training without interactions with the dedicated environments, we introduce a data synthesis technique that autonomously collects conversational datasets by swapping the roles of the agent and environment. This method is especially useful in scenarios where only a predefined prompt is available without a corresponding training dataset. By internalizing complex prompts, Generative Context Distillation enables high-performance and efficient inference without the need for explicit prompts. [^1] Code are available at this link.
@@ 1 Introduction
To address this issue, existing methods can be categorized into two approaches. One practical approach only for efficient prompt computation is compressing prompts into external embeddings (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024) or compressed text  (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). However, these methods still require additional tokens, limiting their ability to fully internalize prompts and effectively modify the model’s behavior as intended. Alternatively, fine-tuning (Kim and Rush, 2016; Zou et al., 2024) or distilling, a.k.a context distillation (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), have been explored as canonical approaches for internalizing prompts. These methods adjust the language model’s behavior to follow the prompt’s intentions without requiring a prompt during inference.
@@
To overcome the limitations, we propose Generative Context Distillation (GCD), a method that is trained to generate the target prompt, rather than merely using it as input. GCD employs joint loss training, combining two approaches: 1) mimicking the teacher’s output to guide behavior, similar to context distillation approaches (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), and 2) generating the content of the prompt while inferring why the output should change based on that prompt. GCD employs a lightweight internalization process for each prompt, requiring only 0.5% additional parameters on 1,000 samples. Focusing on scenario where only prompts are available without a corresponding training dataset for internalization (Liu et al., 2023; Choi et al., 2023), we also introduce a data synthesis technique to generate a multi-turn pseudo conversational dataset. By simply swapping the roles of user and assistant within the given context, the model simulates both sides of conversation, enabling it to autonomously collect conversational datasets.
We evaluate our method in agent application scenarios with lengthy prompts using AgentBench (Liu et al., 2023). GCD maintains strong performance even without the need for prompt input, while outperforming other distillation or compression-based baselines. GCD achieves 100% performance retention on OS interaction agent task, while maintaining at least 82% performance on web-based agent tasks with over 1,000 tokens. Furthermore, GCD demonstrates 39% efficiency improvement in handling environments with lengthy prompts, outperforming other compression-based methods.
@@ 2 Related Works
Context Distillation
is a methodology designed to internalize context tokens within a language model, enabling the model to perform various tasks without requiring explicit contextual information. For instance, Askell et al. (2021) distills a persona as context to facilitate helpful, honest, and harmless alignment, while Snell et al. (2022) incorporates both prompts and scratch pads into the context distillation process to enhance performance on more complex tasks. Li et al. (2024) applies context distillation to internalize demonstrations for in-context learning scenarios. While these methods focus on distilling generalized and coarse-grained contexts, Choi et al. (2023); Zou et al. (2024) introduce a fine-grained context distillation approach, targeting more specific and predetermined prompts.  Choi et al. (2023) focuses on short chat histories or prompts, and Zou et al. (2024) relies on a tailored training dataset to retrieve similar examples based on the given user input. Following these previous works, our goal is to enable the model to internalize the specific predetermined prompts. However, existing methods still limit the model’s ability to directly learn the content of the prompt, as they rely on training a student model using outputs from a teacher model. We address this limitation by generating prompts that allow the model to learn the content directly, enabling it to handle more realistic and information-rich prompts.
@@ 3 Problem Definition
Following Choi et al. (2023), we assume a scenario where an application-specific prompt $p$ is predetermined. Our goal is to guide the model to behave as if the prompt is given, even in its absence. Similar to the previous context distillation approaches (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), we define the teacher model $T$ and the student model $S$ based on whether the prompt $p$ is provided to the same model $\theta$.
@@
4 Generative Context Distillation
Figure 1: Overview of Generative Context Distillation. SFT loss learns the teacher model’s behavior based on the user input. Prompt Generation loss internalizes the prompt by generating both the content of the prompt and the reason for why the model’s behavior should be modified. This process is guided by comparing the student model’s output (“AS-IS”) with the teacher model’s output (“TO-BE”). SFT loss and Prompt Generation loss are combined into a joint loss to train the prompt-specific adaptor.
In this section, we demonstrate our novel method, Generative Context Distillation (GCD), to internalize the given prompt. Basically, GCD involves training the student model to mimic the behavior of the teacher model given the prompt $p$. We finetune the student model on the multi-turn outputs generated by the teacher model ($y^{T}_{i},i\in\{0,1,...,N\}$), inspired by sequence-level knowledge distillation approaches (Kim and Rush, 2016; Touvron et al., 2023b) which follow the hard label distribution from the teacher model.
@@ 6.2 Baselines
Since previous typical context distilled language models are based on much smaller model (Choi et al., 2023) or are not publicly available  (Snell et al., 2022), we adopt recent knowledge distillation approaches Lin et al. (2020); Gu et al. (2024); Ko et al. (2024) for autoregressive large language models in our context distillation setup. (1) The basic distillation approach (Hinton et al., 2015) employs the Kullback-Leibler Divergence loss (KLD) between the logit distributions of the student and teacher model. (2) Sequence-level Knowledge Distillation (SeqKD) (Kim and Rush, 2016) enforces the student model to generate the teacher model’s outputs on a fixed dataset. (3) As a strong baseline, we also employ a joint loss (SeqKD+KLD), inspired by recent knowledge distillation works (Song et al., 2020; Gu et al., 2024), which incorporates the language modeling loss during distillation. This approach can be interpreted as a hybrid distillation loss that combines the benefits of both soft labels and hard labels from the teacher model.
@@ 7.1 Performance
Generative Context Distillation.
GCD achieves superior performance in the OS interaction task, reaching 100% of the upper bound. It also consistently demonstrates high performance of over 82% across Web Browsing and Web Shopping tasks, both of which have longer prompts that exceed 1,000 tokens in length. The incorporation of PG loss into SFT loss shows a notable performance improvement of approximately 25% in web shopping tasks. Furthermore, when compared to Prompt Prepending which is based on SFT loss and QLoRA (Dettmers et al., 2023) adaptor, our approach exhibits higher performance by leveraging PG loss, rather than relying solely on prepending prompts to the input.
@@ 7.2 Analysis
To understand the factors behind the performance improvement of GCD, we perform three types of ablation studies. First, we compare the impact of PG loss and SFT loss on task performance. Second, we investigate the influence of the reason and prompt in PG loss by removing each component during generation. Finally, instead of omitting components, we explore the effect of reordering the input and output elements in PG loss while keeping the overall information level constant.
@@ 7.2.2 Impact of the Reason/Prompt
In Table 2, the prompt has a greater impact on performance than the reason across all tasks. This is interpreted as being due to the fact that the prompt is longer in length than the reason and generally contains more information, thereby exerting a more significant influence on the Prompt Generation loss.
@@ 8 Efficiency
To compare the efficiency between Generative Context Distillation and other baselines that require fixed prompts, we sample an example with more than 5 turns from a Web Shopping task containing over 1200 tokens. [^5] Efficiency is measured on a single NVIDIA A6000 GPU and AMD EPYC 7513 CPU, featuring 32 physical cores. We then compare the MACs (Multiply-Accumulate Operations), FLOPs (Floating Point Operations), and latency to LLaMA-3-8B-Instruct (Dubey et al., 2024) based baselines. All metrics were measured by DeepSpeed Profile [^6] https://www.deepspeed.ai/tutorials/flops-profiler/.
@@
In Figure 4, the prompt caching time executed before the first turn incurs higher computational overhead due to the use of a larger model compared to the compression model. From the second turn onward, although caching previous conversation histories eliminates redundant computations, the computational overhead still increases proportionally to the length of the prompt. Even without considering in caching time, GCD demonstrates a 39% improvement in MACs and FLOPs, along with a 17% improvement in latency. Consequently, despite caching, baselines inevitably require more computational overhead compared to Generative Context Distillation.
@@ 9 Conclusion
In this paper, we propose a novel prompt internalization method, Generative Context Distillation (GCD), which generates both the contents of the prompt and the reasoning behind changes to the model’s output while mimicking the behavior of the teacher model. To address the challenging scenario where only predetermined prompts are available without additional training data, we introduce Self Role-Playing Conversation, a method that generates a pseudo-conversational training dataset from the given prompt. Our approach demonstrates that even in scenarios with lengthy agent prompts, GCD maintains high performance without relying on the prompt itself. Moreover, GCD improves efficiency by fully internalizing the prompt without requiring any additional tokens.
@@
We assume the use of long and fixed prompts from realistic application scenarios. However, collecting official prompts from real-world applications (such as ChatGPT) poses significant challenges, leading us to rely on prompts from academic agent-based applications. This introduces a limitation, as only a representative prompt for each task are considered, and our approach handles a relatively small set of prompts. In future work, we plan to explore context internalization in various domains, including long chat histories, in-context learning, and retrieval-augmented generation (RAG), as well as in multimodal applications involving video and image data.
@@ Acknowledgements
This research was supported by the MSIT (Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program (RS-2024-00436680) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation). And, this project was also supported by Microsoft Research Asia.
@@ C.4 Implementation Details
We utilize a single NVIDIA A6000 GPU to train our method and baselines. For Generative Context Distillation, we apply QLoRA (Dettmers et al., 2023) with rank $r=16$, 32 batch size, 1e-4 learning rate, 10 epochs. The joint loss hyperparameter $\lambda=0.7$ (as defined in Equation 4) is selected through hyperparameter tuning over [0.3, 0.5, 0.7]. For full finetining baselines, we use 4 NVIDIA A100 80G GPUs with 64 batch size, 1e-5 learning rate, 10 epochs, following the setting from Gu et al. (2024). Prompt Prepending, a variation of our approach, uses the same hyperparameters as GCD. For ICAE (Ge et al., 2024), we follow the original settings with a LoRA rank of $r=512$ and 128 compression tokens.
@@ A. None of the above
Figure 7: Agent Prompt for Web Shopping. Following the task setup from AgentBench (Liu et al., 2023), we describe all content, including the system prompt and demonstrations, as a multi-turn strategy using <USER> and <AGENT>.

### v2

# Generative Prompt Internalization

@@ Abstract
Prompts used in recent large language model based applications are often fixed and lengthy, leading to significant computational overhead. To address this challenge, we propose Generative Prompt Internalization (GenPI), a lightweight method that employs a joint training approach. GenPI not only replicates the behavior of models with prompt inputs but also generates the content of the prompt along with reasons for why the model’s behavior should change accordingly. We demonstrate that our approach effectively internalizes complex prompts across various agent-based application scenarios. For effective training without interactions with the dedicated environments, we introduce a data synthesis technique that autonomously collects conversational datasets by swapping the roles of the agent and environment. This method is especially useful in scenarios where only a predefined prompt is available without a corresponding training dataset. By internalizing complex prompts, Generative Prompt Internalization enables high-performance and efficient inference without the need for explicit prompts. [^1] Code are available at https://github.com/kaistai/GenPI.
@@ 1 Introduction
To address this issue, existing methods can be categorized into two approaches. One practical approach only for efficient prompt computation is compressing prompts into external embeddings (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024) or compressed text  (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). However, these methods still require additional tokens, limiting their ability to fully internalize prompts and effectively modify the model’s behavior as intended. Alternatively, fine-tuning (Kim and Rush, 2016; Zou et al., 2024) or distillation (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), have been explored as canonical approaches for internalizing prompts. These methods adjust the language model’s behavior to follow the prompt’s intentions without requiring a prompt during inference.
@@
To overcome the limitations, we propose Generative Prompt Internalization (GenPI), a method that is trained to generate the target prompt, rather than merely using it as input. GenPI employs joint loss training, combining two approaches: 1) mimicking the teacher’s output to guide behavior, similar to distillation approaches (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), and 2) generating the content of the prompt while inferring why the output should change based on that prompt. GenPI employs a lightweight internalization process for each prompt, requiring only 0.5% additional parameters on 1,000 samples. Focusing on scenario where only prompts are available without a corresponding training dataset for internalization (Liu et al., 2023; Choi et al., 2023), we also introduce a data synthesis technique to generate a multi-turn pseudo conversational dataset. By simply swapping the roles of user and assistant within the given context, the model simulates both sides of conversation, enabling it to autonomously collect conversational datasets.
We evaluate our method in agent application scenarios with lengthy prompts using AgentBench (Liu et al., 2023). GenPI maintains strong performance even without the need for prompt input, while outperforming other distillation or compression-based baselines. GenPI achieves 100% performance retention on OS interaction agent task, while maintaining at least 82% performance on web-based agent tasks with over 1,000 tokens. Furthermore, GenPI demonstrates 39% efficiency improvement in handling environments with lengthy prompts, outperforming other compression-based methods.
@@ 2 Related Works
Prompt Internalization
is a methodology designed to embed prompt information within a language model, enabling the model to perform various tasks without requiring explicit prompt input. For instance, Askell et al. (2021) internalize persona-related prompts to facilitate helpful, honest, and harmless alignment, while Snell et al. (2022) incorporates both prompts and scratch pads into the internalization process to enhance performance on more complex tasks. Li et al. (2024) applies prompt internalization to internalize demonstrations for in-context learning scenarios. While these methods focus on internalizing broad and general coarse-grained information, Choi et al. (2023); Zou et al. (2024) introduce a more fine-grained approach, targeting more specific and predetermined prompts.  Choi et al. (2023) focuses on short chat histories or task-specific instructions, and Zou et al. (2024) relies on a tailored training dataset to retrieve similar examples based on the given user input. Following these previous works, our goal is to enable the model to internalize the specific predetermined prompts. However, existing methods still limit the model’s ability to directly learn the content of the prompt, as they rely on training a student model using outputs from a teacher model. We address this limitation by generating prompts that allow the model to learn their content directly, enabling it to handle more realistic and information-rich prompts.
@@ 3 Problem Definition
Following Choi et al. (2023), we assume a scenario where an application-specific prompt $p$ is predetermined. Our goal is to guide the model to behave as if the prompt is given, even in its absence. Similar to the previous distillation approaches (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), we define the teacher model $T$ and the student model $S$ based on whether the prompt $p$ is provided to the same model $\theta$.
@@
4 Generative Prompt Internalization
Figure 1: Overview of Generative Prompt Internalization. SFT loss learns the teacher model’s behavior based on the user input. Prompt Generation loss internalizes the prompt by generating both the content of the prompt and the reason for why the model’s behavior should be modified. This process is guided by comparing the student model’s output (“AS-IS”) with the teacher model’s output (“TO-BE”). SFT loss and Prompt Generation loss are combined into a joint loss to train the prompt-specific adaptor.
In this section, we demonstrate our novel method, Generative Prompt Internalization (GenPI), which internalizes the given prompt by generating its own contents. Basically, GenPI involves training the student model to mimic the behavior of the teacher model given the prompt $p$. We finetune the student model on the multi-turn outputs generated by the teacher model ($y^{T}_{i},i\in\{0,1,...,N\}$), inspired by sequence-level knowledge distillation approaches (Kim and Rush, 2016; Touvron et al., 2023b) which follow the hard label distribution from the teacher model.
@@ 6.2 Baselines
Since previous typical prompt internalization methods are based on much smaller model (Choi et al., 2023) or are not publicly available  (Snell et al., 2022), we adopt recent knowledge distillation approaches Lin et al. (2020); Gu et al. (2024); Ko et al. (2024) for autoregressive large language models in our prompt internalization setup. (1) The basic distillation approach (Hinton et al., 2015) employs the Kullback-Leibler Divergence loss (KLD) between the logit distributions of the student and teacher model. (2) Sequence-level Knowledge Distillation (SeqKD) (Kim and Rush, 2016) enforces the student model to generate the teacher model’s outputs on a fixed dataset. (3) As a strong baseline, we also employ a joint loss (SeqKD+KLD), inspired by recent knowledge distillation works (Song et al., 2020; Gu et al., 2024), which incorporates the language modeling loss during distillation. This approach can be interpreted as a hybrid distillation loss that combines the benefits of both soft labels and hard labels from the teacher model.
@@ 7.1 Performance
Generative Prompt Internalization.
GenPI achieves superior performance in the OS interaction task, reaching 100% of the upper bound. It also consistently demonstrates high performance of over 82% across Web Browsing and Web Shopping tasks, both of which have longer prompts that exceed 1,000 tokens in length. The incorporation of PG loss into SFT loss shows a notable performance improvement of approximately 25% in web shopping tasks. Furthermore, when compared to Prompt Prepending which is based on SFT loss and QLoRA (Dettmers et al., 2023) adaptor, our approach exhibits higher performance by leveraging PG loss, rather than relying solely on prepending prompts to the input.
@@ 7.2 Analysis
To understand the factors behind the performance improvement of GenPI, we perform three types of ablation studies. First, we compare the impact of PG loss and SFT loss on task performance. Second, we investigate the influence of the reason and prompt in PG loss by removing each component during generation. Finally, instead of omitting components, we explore the effect of reordering the input and output elements in PG loss while keeping the overall information level constant.
@@ 7.2.2 Impact of the Reason/Prompt
In Table 2, the prompt $p$ has a greater impact on performance than the reason $r$ across all tasks. This is interpreted as being due to the fact that the prompt is longer in length than the reason and generally contains more information, thereby exerting a more significant influence on the Prompt Generation loss. This highlights our primary contribution, directly generating the prompt itself, allowing the model to internalize and utilize prompt information more effectively.
@@ 8 Efficiency
To compare the efficiency between Generative Prompt Internalization and other baselines that require fixed prompts, we sample an example with more than 5 turns from a Web Shopping task containing over 1200 tokens. [^5] Efficiency is measured on a single NVIDIA A6000 GPU and AMD EPYC 7513 CPU, featuring 32 physical cores. We then compare the MACs (Multiply-Accumulate Operations), FLOPs (Floating Point Operations), and latency to LLaMA-3-8B-Instruct (Dubey et al., 2024) based baselines. All metrics were measured by DeepSpeed Profile [^6] https://www.deepspeed.ai/tutorials/flops-profiler/.
@@
In Figure 4, the prompt caching time executed before the first turn incurs higher computational overhead due to the use of a larger model compared to the compression model. From the second turn onward, although caching previous conversation histories eliminates redundant computations, the computational overhead still increases proportionally to the length of the prompt. Even without considering in caching time, GenPI demonstrates a 39% improvement in MACs and FLOPs, along with a 17% improvement in latency. Consequently, despite caching, baselines inevitably require more computational overhead compared to Generative Prompt Internalization.
9 Discussion
In this section, we discuss the strengths and limitations of two prompt internalization strategies: generalized prompt-agnostic compression and prompt-specific internalization approach.
Prompt compression approaches, such as LLMLingua-2 (Pan et al., 2024) and ICAE (Ge et al., 2024), effectively compress short instructions without requiring additional training for new prompts (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024). However, as shown in Table 1, these methods struggle with long, information-rich application prompts, where key details are often lost in the compression process.
Prompt internalization approach, such as GenPI, overcome this limitation by explicitly internalizing prompts through targeted training. While GenPI requires training for each new prompt, this overhead is minimal in practical applications where only a limited set of prompts is needed. Once trained, switching between adapters incurs negligible computational cost, making the approach both flexible and scalable. Each adapter requires only 1,000 data samples for training. Even when generating the dataset with a 70B LLM, entire process can be completed in under an hour using four A6000 GPUs with vLLM (Kwon et al., 2023), while training takes approximately three hours on a single A6000 GPU.
@@ 10 Conclusion
In this paper, we propose a novel prompt internalization method, Generative Prompt Internalization (GenPI), which generates both the contents of the prompt and the reasoning behind changes to the model’s output while mimicking the behavior of the teacher model. To address the challenging scenario where only predetermined prompts are available without additional training data, we introduce Self Role-Playing Conversation, a method that generates a pseudo-conversational training dataset from the given prompt. Our approach demonstrates that even in scenarios with lengthy agent prompts, GenPI maintains high performance without relying on the prompt itself. Moreover, GenPI improves efficiency by fully internalizing the prompt without requiring any additional tokens.
@@
We assume the use of long and fixed prompts from realistic application scenarios. However, collecting official prompts from real-world applications (such as ChatGPT) poses significant challenges, leading us to rely on prompts from academic agent-based applications. This introduces a limitation, as only a representative prompt for each task are considered, and our approach handles a relatively small set of prompts. In future work, we plan to further explore context internalization in various domains, including long chat histories, in-context learning, and retrieval-augmented generation (RAG), as well as multimodal applications involving video and image data.
@@ Acknowledgements
This project was supported by Microsoft Research Asia. Additionally, this research was supported by the MSIT (Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program (RS-2024-00436680) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation). And, this work also supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (RS-2024-00398115, Research on the reliability and coherence of outcomes produced by Generative AI).
@@ C.4 Implementation Details
We utilize a single NVIDIA A6000 GPU to train our method and baselines. For Generative Prompt Internalization, we apply QLoRA (Dettmers et al., 2023) with rank $r=16$, 32 batch size, 1e-4 learning rate, 10 epochs. The joint loss hyperparameter $\lambda=0.7$ (as defined in Equation 4) is selected through hyperparameter tuning over [0.3, 0.5, 0.7]. For full finetining baselines, we use 4 NVIDIA A100 80G GPUs with 64 batch size, 1e-5 learning rate, 10 epochs, following the setting from Gu et al. (2024). Prompt Prepending, a variation of our approach, uses the same hyperparameters as GenPI. For ICAE (Ge et al., 2024), we follow the original settings with a LoRA rank of $r=512$ and 128 compression tokens.
Appendix D Comparison with Chain-of-Thought
To discuss of the role of reason $r$ in PG loss, we compare the differences between chain-of-thought (CoT) rationales (Wei et al., 2022) and reason $r$, focusing on their mechanisms, and impact on inference.
Rationale vs. Behavioral Change.
CoT (Wei et al., 2022) generates explicit stepwise rationales to explain the final answer for solving a task-intensive problem (e.g., an explicit reasoning process for solving a math problem). In contrast, reason $r$ aims to compare the differences between the outputs of the teacher and student models, and identify the key changes in the model’s behavior. That is, while indirectly describing the prompt $p$, PG loss is designed to induce behavioral changes in the model rather than explicitly providing explanations for the correct answer.
Inference Efficiency.
Since GenPI does not generate rationales at inference time, it differs from CoT (Wei et al., 2022) in inference efficiency. By internalizing the prompt, PG loss ensures higher efficiency (in Section 8) while maintaining the original performance.
@@ A. None of the above
Figure 7: Agent Prompt for Web Shopping. Following AgentBench (Liu et al., 2023), we describe all content, including the system prompt and demonstrations, as a multi-turn strategy using <USER> and <AGENT>.